Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by raunakrajan » Wed Jun 30, 2010 3:21 am
A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)

[spoiler]1/2[/spoiler]
Join the discussion
Source: — Problem Solving |

by kvcpk » Wed Jun 30, 2010 3:59 am
raunakrajan wrote:A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)

[spoiler]1/2[/spoiler]
1st display: 2r 2b
So probability of choosing a r is 2/4 and probability of choosing a b is 2/4

2nd display: 3r 2b
So probability of choosing a r is 3/5 and probability of choosing a b is 2/5

we need 1 red and 1 blue to get violet.

Prob that r is chosen from 1st disp and b from 2nd disp is
2/4 * 2/5

Prob that b is chosen from 1st disp and r from 2nd disp is
2/4 * 3/5

We need either of these two to happen.

So 2/4 * 2/5 +2/4 * 3/5
= 2/4 = 1/2

Hope this helps!!
Join the discussion

by raunakrajan » Wed Jun 30, 2010 4:02 am
kvcpk wrote:
raunakrajan wrote:A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)

[spoiler]1/2[/spoiler]
1st display: 2r 2b
So probability of choosing a r is 2/4 and probability of choosing a b is 2/4

2nd display: 3r 2b
So probability of choosing a r is 3/5 and probability of choosing a b is 2/5

we need 1 red and 1 blue to get violet.

Prob that r is chosen from 1st disp and b from 2nd disp is
2/4 * 2/5

Prob that b is chosen from 1st disp and r from 2nd disp is
2/4 * 3/5

We need either of these two to happen.

So 2/4 * 2/5 +2/4 * 3/5
= 2/4 = 1/2

Hope this helps!!
thanks a lot!!
Join the discussion

by kvcpk » Wed Jun 30, 2010 4:13 am
You are welcome!! :)
Join the discussion

probability

by GMATGuruNY » Wed Jun 30, 2010 5:47 am
raunakrajan wrote:A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)
Case 1: P(R then B) = 2/4 * 2/5 = 4/20.
Case 2: P(B then R) = 2/4 * 3/5 = 6/20.
Since a good outcome will be yielded by Case 1 OR Case 2, we ADD the probabilities:
4/20 + 6/20 = 1/2.
Last edited by GMATGuruNY on Sun Jun 14, 2020 3:31 am, edited 1 time in total.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by kvcpk » Wed Jun 30, 2010 7:40 am
GMATGuruNY wrote:
raunakrajan wrote:A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)

[spoiler]1/2[/spoiler]
This question could be rephrased:

What is the probability that the painter picks the right color when he chooses the second jar?

The probability of picking R or B on any given pick will be the same as the probability of picking R or B on the first pick. (See https://www.beatthegmat.com/probablity-ques-t60161.html for a discussion of this concept.)

Since the painter has an equal chance of gettting R or B on his first pick, he has a 1/2 chance of getting the right color on his 2nd pick as well.

The correct answer is 1/2.
I think it will take tonnes of practice to get this level of thinking.. Thanks for sharing!!
Join the discussion

by GMATGuruNY » Wed Jun 30, 2010 12:22 pm
kvcpk wrote:
GMATGuruNY wrote:
raunakrajan wrote:A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)

[spoiler]1/2[/spoiler]
This question could be rephrased:

What is the probability that the painter picks the right color when he chooses the second jar?

The probability of picking R or B on any given pick will be the same as the probability of picking R or B on the first pick. (See https://www.beatthegmat.com/probablity-ques-t60161.html for a discussion of this concept.)

Since the painter has an equal chance of gettting R or B on his first pick, he has a 1/2 chance of getting the right color on his 2nd pick as well.

The correct answer is 1/2.
I think it will take tonnes of practice to get this level of thinking.. Thanks for sharing!!
I agree. The streamlined approach that I took is efficient and slick, but it might not be immediately apparent to every test-taker. But the concept that was applied is worth noting:

If a question asks for the probability of getting X on the nth pick, the answer will be the probability of getting X on the FIRST pick.

So if you have 100 green M&Ms and 80 red M&Ms in a huge jar, and you select 50 because you have an insatiable sweet tooth, and the question asks:

What is the probability that the 43rd M&M selected is green?

The answer will be the probability of selecting a green M&M on your FIRST pick: 100/180 = 10/18 = 5/9.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by raunakrajan » Wed Jun 30, 2010 12:28 pm
GMATGuruNY wrote:
kvcpk wrote:
GMATGuruNY wrote:
raunakrajan wrote:A painter intends to mixed two jars of paint to create a new color. He chooses one paint jar randomly from a display containing 2 red and 2 blue samples, and a second jar is chosen at random from a different display containing 3 red and 2 blue samples. If he plans on mixing the two chosen jars together, what is the probability that when mixed together, the result will be purple paint? (The color purple is created with 1 part red and 1 part blue.)

[spoiler]1/2[/spoiler]
This question could be rephrased:

What is the probability that the painter picks the right color when he chooses the second jar?

The probability of picking R or B on any given pick will be the same as the probability of picking R or B on the first pick. (See https://www.beatthegmat.com/probablity-ques-t60161.html for a discussion of this concept.)

Since the painter has an equal chance of gettting R or B on his first pick, he has a 1/2 chance of getting the right color on his 2nd pick as well.

The correct answer is 1/2.
I think it will take tonnes of practice to get this level of thinking.. Thanks for sharing!!
I agree. The streamlined approach that I took is efficient and slick, but it might not be immediately apparent to every test-taker. But the concept that was applied is worth noting:

If a question asks for the probability of getting X on the nth pick, the answer will be the probability of getting X on the FIRST pick.

So if you have 100 green M&Ms and 80 red M&Ms in a huge jar, and you select 50 because you have an insatiable sweet tooth, and the question asks:

What is the probability that the 43rd M&M selected is green?

The answer will be the probability of selecting a green M&M on your FIRST pick: 100/180 = 10/18 = 5/9.
thanks Mitch!
if you could explain the same sum (M&M's) in the traditional way how do we go about it!
to be honest I couldn't understand the logic
Join the discussion

by GMATGuruNY » Wed Jun 30, 2010 12:59 pm
If a question asks for the probability of getting X on the nth pick, the answer will be the probability of getting X on the FIRST pick.

Let's prove this theory by using easier numbers. Say we have 3 green M&Ms and 4 red M&Ms, and we select 3 at random.

What's the probability that the first M&M selected will be green?

3/7

What's the probability that the second M&M selected will be green?

Two ways to get a good outcome. We could select (G and G) or (R and G):

P(G and G) = 3/7 * 2/6 = 1/7.
P(R and G) = 4/7 * 3/6 = 2/7.

Since either is a good outcome, we add the fractions: 1/7 + 2/7 = 3/7.

What's the probability that the third M&M selected will be green?

Several ways to get a good outcome. We could select (G and G and G), (G and R and G), R and G and G), or (R and R and G):

P(G and G and G) = 3/7 * 2/6 * 1/5 = 1/35
P(G and R and G) = 3/7 * 4/6 * 2/5 = 4/35
P(R and G and G) = 4/7 * 3/6 * 2/5 = 4/35
P(R and R and G) = 4/7 * 3/6 * 3/5 = 6/35

Since any of the above would give us a good outcome, we add the fractions:

1/35 + 4/35 + 4/35 + 6/35 = 15/35 = 3/7.

Do you see what's happening?

The probability that the first M&M selected is green = 3/7.
The probability that the second M&M selected is green = 3/7.
The probability that the third M&M selected is green = 3/7.

We keep getting the same fraction: 3/7. The probability that the third M&M selected is green = the probability that the second M&M selected is green = the probability that the first M&M selected is green.

So as noted above:

If a question asks for the probability of getting X on the nth pick, the answer will be the probability of getting X on the FIRST pick.

This rule holds true even if we increase the numbers. So if you have 100 green M&Ms and 80 red M&Ms in a huge jar, and you select 50 because you have an insatiable sweet tooth, and the question asks:

What is the probability that the 43rd M&M selected is green?

The answer will be the probability of selecting a green M&M on your FIRST pick: 100/180 = 10/18 = 5/9.

Does this help?
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by raunakrajan » Wed Jun 30, 2010 1:03 pm
GMATGuruNY wrote:If a question asks for the probability of getting X on the nth pick, the answer will be the probability of getting X on the FIRST pick.

Let's prove this theory by using easier numbers. Say we have 3 green M&Ms and 4 red M&Ms, and we select 3 at random.

What's the probability that the first M&M selected will be green?

3/7

What's the probability that the second M&M selected will be green?

Two ways to get a good outcome. We could select (G and G) or (R and G):

P(G and G) = 3/7 * 2/6 = 1/7.
P(R and G) = 4/7 * 3/6 = 2/7.

Since either is a good outcome, we add the fractions: 1/7 + 2/7 = 3/7.

What's the probability that the third M&M selected will be green?

Several ways to get a good outcome. We could select (G and G and G), (G and R and G), R and G and G), or (R and R and G):

P(G and G and G) = 3/7 * 2/6 * 1/5 = 1/35
P(G and R and G) = 3/7 * 4/6 * 2/5 = 4/35
P(R and G and G) = 4/7 * 3/6 * 2/5 = 4/35
P(R and R and G) = 4/7 * 3/6 * 3/5 = 6/35

Since any of the above would give us a good outcome, we add the fractions:

1/35 + 4/35 + 4/35 + 6/35 = 15/35 = 3/7.

Do you see what's happening?

The probability that the first M&M selected is green = 3/7.
The probability that the second M&M selected is green = 3/7.
The probability that the third M&M selected is green = 3/7.

We keep getting the same fraction: 3/7. The probability that the third M&M selected is green = the probability that the second M&M selected is green = the probability that the first M&M selected is green.

So as noted above:

If a question asks for the probability of getting X on the nth pick, the answer will be the probability of getting X on the FIRST pick.

This rule holds true even if we increase the numbers. So if you have 100 green M&Ms and 80 red M&Ms in a huge jar, and you select 50 because you have an insatiable sweet tooth, and the question asks:

What is the probability that the 43rd M&M selected is green?

The answer will be the probability of selecting a green M&M on your FIRST pick: 100/180 = 10/18 = 5/9.

Does this help?
Bingo!
thank you so much Mitch!
this seems so doable now!
cant thank you enough!

Best
Raunak
Join the discussion